ab-5=b/3
Solve for b

Answer by stanbon(48568)

ab-5=b/3
Solve for b
-------------
Multiply thru by 3 to get:
3ab-15 = b
Get the "b's" on one side:
3ab-b = 15
Factor out the "b":
b(3a-1) = 15
Divide to solve for "b".
b = 15/(3a-1)
===========
Cheers,
Stan H.